# thanks: $0$-0002-5209-872Xthanks: $0$-0002-1235-0505

### v1

@@ III Sensitivities
Our results show that CID can improve the precision on $\delta$ by 15-20%. We have also studied the effect of assuming muon and electron CID at the same time and the results are similar. We find that electron CID has a bigger impact on the improvements in precision on $\delta$. This is likely due to the fact that the detected $\nu_{\mu}\to\nu_{e}$ spectrum is more similar to the detected $\bar{\nu}_{e}\to\bar{\nu}_{e}$ spectrum than the detected $\nu_{e}\to\nu_{\mu}$ spectrum to the detected $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ spectrum, see appendix A. We find that the sensitivity of a neutrino factory comes somewhat more from the $\nu_{e}\to\nu_{\mu}$ golden channel De Rujula et al. (1999); Cervera et al. (2000) and less from the $\nu_{\mu}\to\nu_{e}$ channel which will be leveraged by future proton accelerator neutrino experiments. The oscillation probabilities in these two channels are the same up to the change $\delta\to-\delta$, so each channel serves as an important cross check of the other, but provides no fundamentally new insights into the nature of CP violation or conservation. They do, however, provide a probe of CPT invariance because a neutrino factory is sensitive to four different appearance channels which CP, T, and CPT conjugates of each other. Thus a neutrino factory, especially when combined with DUNE and HK, will provide a valuable unique test of CPT invariance.
@@
Overall, the improvements on precision on $\delta$ with the BNL-SURF setup are larger than at the FNAL-SURF setup largely because in the former setup the neutrinos experience more matter effects Wolfenstein (1978a) due to the higher neutrino energy which increases the appearance probability.
@@
While increased precision on $\delta$ and other oscillation parameters for the sake of precision is an important goal of particle physics, we also provide some context for these sensitivities. Because $\delta$ is the last remaining oscillation parameter that is largely unconstrained, it plays a key role in flavor model predictions Girardi et al. (2015); Everett et al. (2021), along with other parameters such as $\theta_{12}$ and $\Delta m^{2}_{21}$ Gehrlein et al. (2022) which are expected to become measured with much more precision in the coming years by JUNO Abusleme et al. (2022). While there are many conceivable predictions on neutrino parameters from many different structures (see Denton and Gehrlein (2024) for a recent survey), here we focus on charged lepton corrections as they predict the mixing parameters, but not the masses. In fig. 3 we show the predictions of $\cos\delta$ from four popular charged lepton correction structures as discussed in Everett et al. (2021) alongside the expected $1\sigma$ precision on $\cos\delta$ (taking the wider allowed ranges in all cases due to the sign degeneracy since long-baseline experiments predominantly measure $\sin\delta$ from their appearance channels, although they also probe $\cos\delta$ from their disappearance measurements Denton (2023)). We consider both the precision from the combination of DUNE and HK as well as the expected improvement from including the neutrino factory at the FNAL baseline. In table 1 we document the expected precision on $\delta$ for various experiments.
@@
Figure 3: The predicted regions of $\cos\delta$ due to several popular structures within the charged-lepton correction framework, adapted from Everett et al. (2021); GR1 and GR2 are two flavor model predictions based on the golden ratio Datta et al. (2003); Everett and Stuart (2009); Rodejohann (2009); Adulpravitchai et al. (2009); Feruglio and Paris (2011); Ding et al. (2012), TBM is the tribimaximal flavor model prediction Wolfenstein (1978b); Harrison et al. (2002); Harrison and Scott (2002); Xing et al. (2006); He and Zee (2003), and HEX is based on a hexagonal structure Albright et al. (2010). Along the top are the expected $1\sigma$ sensitivities to $\cos\delta$ at the various central values predicted by the models. DUNE and HK are both at 5+5 years as described in the text, as is neutrino factory with the FNAL baseline.
@@
We see that while after DUNE and HK there will be some modest discriminating capabilities in some cases, the addition of neutrino factory will considerably increase the capability to differentiate flavor models. Combined with JUNO’s constraint on solar parameters should lead to further discrimination capabilities as well.
IV Discussion and Recommendations
@@
Our results show that CID plays a smaller role than emphasized in studies a decade ago. This is primarily due to the fact that LArTPC detectors have sufficiently good energy resolution to distinguish signal from background as well as identify the different flavors. In fact, the LArTPC energy resolution could be even better than assumed here, see e.g. De Romeri et al. (2016); Friedland and Li (2019).
@@
Using a similar detector setup for a neutrino factory as for near-future experiments could further reduce the uncertainties on the cross section and detector efficiencies, as there is an ongoing program to measure the neutrino cross section on Ar and advancing the LAr technology Machado et al. (2019).
A neutrino factory could be an appealing possible future neutrino oscillation experiment should the results of HK and DUNE disagree. If the tension between NOvA and T2K in the CP violation measurements Denton et al. (2021); Chatterjee and Palazzo (2021) persists, DUNE will be able to probe it Denton et al. (2023). However if the tension requires further oscillation studies at a higher neutrino energy, longer baseline, and overall smaller flux uncertainties than fixed target oscillation experiments a neutrino factory is a favorable setup for future studies. [^1] Note that atmospheric neutrino experiments can also provide oscillation studies at higher energies and longer baselines however their flux uncertainty and flavor composition is more uncertain than at a neutrino factory. Furthermore, at a neutrino factory the neutrino energy is flexible and tunable Delahaye et al. (2013). Finally, at a neutrino factory six oscillation channels and their CP conjugate ones are accessible with similar large number of events as the initial neutrino beam consists equally of $\nu_{e},\nu_{\mu}$.
Table 1: Resolution on $\delta$ for combinations of different experiments. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume 10 years for a 40 kT LAr detector and $10^{21}$ muon decays per year.
@@
DUNE $(17.7^{\circ},9.3^{\circ})$
@@
DUNE+HK+NF(FNAL) $(10.6^{\circ},3.8^{\circ})$ $(8.3^{\circ},3.1^{\circ})$ $(8.9^{\circ},3.3^{\circ})$
DUNE+HK+NF(BNL) $(9.2^{\circ},3.9^{\circ})$ $(7.9^{\circ},3.3^{\circ})$ $(8.5^{\circ},3.4^{\circ})$
V Conclusions
Motivated by the recent P5 report, we study the physics potential of a neutrino factory in improving the precision on the determination of the remaining undetermined neutrino oscillation parameters with a focus on the complex phase $\delta$ which governs the amount of CP violation in the neutrino sector. We consider two different baselines and muon energies with a LAr detector based at SURF. Similar baselines and energies were considered as a low-energy neutrino factory Geer et al. (2007); Bross et al. (2008) appropriate to measure CPV for large values of $\theta_{13}$ Tang and Winter (2010); Dighe et al. (2012); Ballett and Pascoli (2012); Christensen et al. (2013). Such a neutrino factory setup could be achieved with a proton beam with 2-4 MW Hol (2000); Oza (2001); Albright et al. (2004); Geer and Zisman (2007); Ankenbrandt et al. (2009).
@@ V Conclusions
We also investigated the role of charge identification of electrons or muons (or both). We found that CID can lead to some improvement in the sensitivity to CP violation. The impact is largest with larger statistics as CID helps to cut through systematics which become relevant at high statistics.
@@ Acknowledgements.
PBD acknowledges support by the United States Department of Energy under Grant Contract No. DE-SC0012704.
@@ Appendix A Number of events
Figure 4: Number of events at a neutrino factory and DUNE for the BNL-SURF and FNAL-SURF baseline with muon energy 8 GeV and 5 GeV for 5+5 years of running. We show the number of events from the $\nu_{e}\to\nu_{\mu}$ channel in red for varying values of $\delta$ and the dominant background from $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$.
@@
Appendix B Results for BNL baseline
In fig. 5 we show the resolution of $\delta=0,-90^{\circ}$ for the baseline between BNL-SURF. We see that a BNL based neutrino factory would lead to significant improvement in the precision on $\delta$, especially for $\delta=-90^{\circ}$, with only a modest benefit from CID.
Figure 5: Resolution on $\delta$ as a function of exposure of a neutrino factory with baseline 2542.3 km, corresponding to the distance between BNL-SURF combined with DUNE and HK. We also show the results for 100% eCID and 100% $\mu$CID as well as without CID.

### v2

@@ 3.1 Sensitivity to $\delta$
We find that the sensitivity of a neutrino factory comes somewhat more from the $\nu_{e}\to\nu_{\mu}$ golden channel DeRujula:1998umv ; Cervera:2000kp and less from the $\nu_{\mu}\to\nu_{e}$ channel which will be leveraged by future proton accelerator neutrino experiments. The oscillation probabilities in these two channels are the same up to the change $\delta\to-\delta$, so each channel serves as an important cross check of the other, but provides no fundamentally new insights into the nature of CP violation or conservation. They do, however, provide a probe of CPT invariance because a neutrino factory is sensitive to four different appearance channels which CP, T, and CPT conjugates of each other. Thus a neutrino factory, especially when combined with DUNE and HK, will provide a valuable unique test of CPT invariance.
@@
Overall, the improvements on precision on $\delta$ with the BNL-SURF setup are larger than at the FNAL-SURF setup largely because in the former setup the neutrinos experience more matter effects Wolfenstein:1977ue due to the higher neutrino energy which increases the appearance probability. Furthermore, due to the higher neutrino energy in this setup the cross section uncertainty is smaller. Finally, even though the flux is lower in the BNL-SURF setup due to the larger baseline and beam spreading, the cross section is higher, the appearance probability is somewhat larger, and the variation in the probability due to variations in $\delta$ is also somewhat larger; these effects provide some cancellation for the lower flux.
@@
While increased precision on $\delta$ and other oscillation parameters for the sake of precision is an important goal of particle physics, we also provide some context for these sensitivities. Because $\delta$ is the last remaining oscillation parameter that is largely unconstrained, it plays a key role in flavor model predictions Girardi:2014faa ; Everett:2019idp , along with other parameters such as $\theta_{12}$ and $\Delta m^{2}_{21}$ Gehrlein:2022nss which are expected to become measured with much more precision in the coming years by JUNO JUNO:2022mxj . While there are many conceivable predictions on neutrino parameters from many different structures (see Denton:2023hkx for a recent survey), here we focus on charged lepton corrections as they predict the mixing parameters and not the masses. In fig. 5 we show the predictions of $\cos\delta$ from four popular charged lepton correction structures as discussed in Everett:2019idp alongside the expected $1\sigma$ precision on $\cos\delta$ (taking the wider allowed ranges in all cases due to the sign degeneracy since long-baseline experiments predominantly measure $\sin\delta$ from their appearance channels, although they also probe $\cos\delta$ from their disappearance measurements Denton:2023qmd ). We consider both the precision from the combination of DUNE and HK as well as the expected improvement from including the neutrino factory at the FNAL baseline. In table 1 we document the expected precision on $\delta$ for various experiments.
@@
Figure 5: The predicted regions of $\cos\delta$ due to several popular structures within the charged-lepton correction framework, adapted from Everett:2019idp ; GR1 and GR2 are two flavor model predictions based on the golden ratio Datta:2003qg ; Everett:2008et ; Rodejohann:2008ir ; Adulpravitchai:2009bg ; Feruglio:2011qq ; Ding:2011cm , TBM is the tribimaximal flavor model prediction Wolfenstein:1978uw ; Harrison:2002er ; Harrison:2002kp ; Xing:2006xa ; He:2003rm , and HEX is based on a hexagonal structure Albright:2010ap . Along the top are the expected $1\sigma$ sensitivities to $\cos\delta$ at the various central values predicted by the models. DUNE and HK are both at 5+5 years as described in the text, as is neutrino factory with the FNAL baseline and no CID.
@@
We see that while after DUNE and HK there will be some modest model discriminating capabilities in some cases, the addition of neutrino factory will increase the capability to differentiate flavor models in some key cases. Significant additional improvement in model discrimination capability is likely to also come from JUNO’s expected improved determination of the solar parameters in coming years as well.
3.2 Sensitivity to other oscillation parameters
In fig. 6 we show the precision on $\sin^{2}\theta_{23},~{}\sin^{2}2\theta_{13}$ and $\Delta m_{31}^{2}$ [^7] Just like DUNE and HK a NF has only very limited sensitivites to the solar parameters Denton:2023zwa . assuming 10 years of a NF with $10^{21}$ muon decays per year combined with 10 years of DUNE and HK each. We compare the results to existing constraints and the expected sensitivies from DUNE and HK alone, and 20 years of DUNE and 10 years of HK. We summarize the results in tab. 2. In fig. 7 we show the correlations between $\sin^{2}\theta_{23}-\Delta m_{31}^{2}$ and $\theta_{13}-\delta$.
We find that the combination of DUNE and HK will drastically improve the precision on the atmospheric parameter $\Delta m_{31}^{2}$ and $\theta_{23}$. The addition of a NF will further improve the precision nearly independent on the assumption of CID. The improvement on the precision on $\Delta m_{31}^{2}$ is slightly larger for the BNL setup where it can also improve over 20 years of DUNE combined with 10 years of HK. For $\theta_{23}$ the improvements in precision due to the addition of a NF are most noticeable for $\sin^{2}\theta_{23}^{\text{true}}=0.5$ where a NF leads to a more symmetric precision. Again we find that the BNL setup leads to slightly higher precision that the FNAL setup and can even improve over 20 years of DUNE plus 10 years of HK. Finally, DUNE and HK lead to slight improvement on the precision of $\sin^{2}2\theta_{13}$ which can even be further increased by the addition of a NF. Both setups could even improve over 20 years of DUNE and 10 years of HK with additional improvements coming from CID of either electrons or muons. Also the two-dimensional precision on $\Delta m_{31}^{2}-\sin^{2}\theta_{23}$ and $\theta_{13}-\delta$ benefit from the addition of a NF to DUNE and HK.
Going to higher muon energies for the BNL setup slightly worsens the oscillation parameters but still provides an improvement over the combination of just DUNE and HK, for some oscillation parameters even compared an increased exposure for DUNE, see tab. 3 in app. B.
Figure 6: Precision on $\Delta m_{31}^{2},~{}\sin^{2}(2\theta_{13}),~{}\sin^{2}\theta_{23}$ for true NO. For $\theta_{23}$ we assume three different true values $\theta_{23}^{\text{true}}=(42^{\circ},~{}45^{\circ},~{}49^{\circ})$. We show the results for 10 years of NF with $10^{21}$ muon decays per year and a 40 kT FD and $E_{\mu}=5~{}(8)$ GeV for the FNAL (BNL) setup on the left (right) plots, and for different assumptions on CID. The results for the NF also include 10 years of both HK and DUNE. We also compare the results to 20 years of DUNE in addition to 10 years of HK. The gray curves show the current precision on these parameters from Esteban:2020cvm ; DayaBay:2022orm .
Figure 7: Two-dimensional precision on the oscillation parameters assuming DUNE+HK+NF (10 years of each) with the FNAL-SURF baseline assuming different true values marked with a star. The contours are drawn at 90% C.L. for 2 dofs.
4 Discussion and Recommendations
This study advances beyond others in several important ways. First, being the first such study in many years, it makes use of improved knowledge of the oscillation parameters; not just $\theta_{13}$, but also $|\Delta m^{2}_{31}|$ which has steadily improved with measurements from 7+ different experiments. Second, the next generation global long-baseline experimental picture is shaping up. There will be a precision reactor experiment (JUNO), two advanced long-baseline accelerator experiments (DUNE and HK), and improvements with atmospheric neutrinos (IceCube, HK, and KM3NeT). Third, the US has committed to a program of LArTPC’s for good energy resolution and particle identification. Fourth, the US interest in a potential muon collider has recently increased leading to the serious possibility of building a muon accelerator, technology that may well require a lower energy machine to demonstrate muon cooling. Finally, we have carefully investigated all relevant aspects of three-flavor oscillations including CID, the role of systematics, and the impact to flavor model predictions.
We now highlight the differences between a neutrino factory and DUNE as presented here. First, the shape of the spectrum is quite different and it tends to rise to the high energy part of the spectrum. This is compared to the broader spectrum that DUNE will see based on the specific horn configuration designed to have good sensitivity to $\delta$ DUNE:2020ypp ; Calviani:2014cxa . Second, the spectrum of neutrinos for a neutrino factory is precisely predicted, while that for DUNE carries significant uncertainties. Third, a neutrino factory provides access to more channels (notably $\nu_{e}\to\nu_{e}$, $\nu_{e}\to\nu_{\mu}$, and their conjugates) than are regularly easily accessible. Fourth, for a given accelerator complex, the flux of neutrinos from a neutrino factory will be similar to, but likely less, than that for DUNE. The exact details here depend on muon cooling efficiencies and the geometry of the neutrino factory, but all factors in this point are no more than $\mathcal{O}(1)$ factors.
@@ 4 Discussion and Recommendations
Our results show that while CID does enhance the oscillation parameters’ precision, CID plays a smaller role than emphasized in studies a decade ago. This is primarily due to the fact that LArTPC detectors have sufficiently good energy resolution to distinguish signal from background as well as identify the different flavors. In fact, the LArTPC energy resolution could be even better than assumed here, see e.g. DeRomeri:2016qwo ; Friedland:2018vry ; Kopp:2024lch .
@@
Using a similar detector setup for a neutrino factory as for near-future experiments could further reduce the uncertainties on the cross section and detector efficiencies, as there is an ongoing program to measure the neutrino cross section on Ar and advancing the LAr technology Machado:2019oxb .
A neutrino factory could be an appealing possible future neutrino oscillation experiment should the results of HK and DUNE disagree. If the weak tension between NOvA and T2K in the CP violation measurements Denton:2020uda ; Chatterjee:2020kkm persists, DUNE will be able to probe it Denton:2022pxt . However if the tension requires further oscillation studies at a higher neutrino energy, longer baseline, and overall smaller flux uncertainties than fixed target oscillation experiments a neutrino factory is a favorable setup for future studies. [^8] Note that atmospheric neutrino experiments can also provide oscillation studies at higher energies and longer baselines however their flux uncertainty and flavor composition is more uncertain than at a neutrino factory. Furthermore, at a neutrino factory the neutrino energy is flexible and tunable Delahaye:2013jla . Finally, at a neutrino factory six oscillation channels and their CP conjugate ones are accessible with similarly large number of events as the initial neutrino beam consists equally of $\nu_{e},\nu_{\mu}$. The presence of numerous oscillation channels provides multiple independent means of getting at the oscillation parameters, each with different dependences on the systematic uncertainties.
Table 1: Resolution on $\delta$ for combinations of different experiments for two true values of $\delta$. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays – nominally 10 years of each experiment unless otherwise specified. We also show the sensitivities for 960 kT-MW-yr of DUNE, twice its nominal exposure, combined with 2.47 MT-MW-yr of HK.
@@
DUNE $(17.8^{\circ},9.4^{\circ})$
@@
DUNE (20 yr)+HK $(11.0^{\circ},4.5^{\circ})$
DUNE+HK+NF(FNAL) $(11.2^{\circ},3.9^{\circ})$ $(8.5^{\circ},3.2^{\circ})$ $(9.0^{\circ},3.3^{\circ})$
DUNE+HK+NF(BNL) $(9.3^{\circ},3.9^{\circ})$ $(8.0^{\circ},3.3^{\circ})$ $(8.6^{\circ},3.4^{\circ})$
Table 2: Resolution on the other parameters for combinations of different experiments. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays – nominally 10 years of each experiment unless otherwise specified. We also show the sensitivities for 960 kT-MW-yr of DUNE, twice its nominal exposure, combined with 2.47 MT-MW-yr of HK.
$\theta_{23}^{\text{true}}=(42^{\circ},~{}45^{\circ},~{}49^{\circ})$ no CID 100% eCID 100% $\mu$CID
HK $(0.22^{\circ},1.37^{\circ},0.23^{\circ})$
DUNE $(0.33^{\circ},1.32^{\circ}0.38^{\circ})$
DUNE+HK $(0.18^{\circ},1.04^{\circ},0.20^{\circ})$
DUNE (20 yr)+HK $(0.15^{\circ},0.92^{\circ},0.16^{\circ})$
DUNE+HK+NF(FNAL) $(0.16^{\circ},0.80^{\circ},0.18^{\circ})$ $(0.16^{\circ},0.74^{\circ},0.18^{\circ})$ $(0.15^{\circ},0.70^{\circ},0.17^{\circ})$
DUNE+HK+NF(BNL) $(0.16^{\circ},0.56^{\circ},0.17^{\circ})$ $(0.15^{\circ},0.52^{\circ},0.17^{\circ})$ $(0.14^{\circ},0.52^{\circ},0.16^{\circ})$
$\theta_{13}^{\text{true}}=8.54^{\circ}$
HK $0.22^{\circ}$
DUNE $0.21^{\circ}$
DUNE+HK $0.16^{\circ}$
DUNE (20 yr)+HK $0.15^{\circ}$
DUNE+HK+NF(FNAL) $0.13^{\circ}$ $0.11^{\circ}$ $0.11^{\circ}$
DUNE+HK+NF(BNL) $0.13^{\circ}$ $0.11^{\circ}$ $0.11^{\circ}$
$(\Delta m_{31}^{2})^{\text{true}}=2.511\cdot 10^{-3}~{}\text{eV}^{2}$ $[10^{-6}~{}\text{eV}^{2}]$ $[10^{-6}~{}\text{eV}^{2}]$ $[10^{-6}~{}\text{eV}^{2}]$
HK $9.8$
DUNE $10.0$
DUNE+HK $7.2$
DUNE (20 yr)+HK $6.1$
DUNE+HK+NF(FNAL) $6.0$ $5.8$ $6.0$
DUNE+HK+NF(BNL) $5.6$ $5.5$ $5.3$
5 Conclusions
Motivated by the recent P5 report, we study the physics potential of a neutrino factory in improving the precision on the determination of the remaining undetermined neutrino oscillation parameters with a focus on the complex phase $\delta$ which governs the amount of CP violation in the neutrino sector. We consider two different baselines and muon energies with a LAr detector based at SURF. Similar baselines and energies were considered as a low-energy neutrino factory Geer:2007kn ; Bross:2007ts appropriate to measure CPV for large values of $\theta_{13}$ Tang:2009wp ; Dighe:2011pa ; Ballett:2012rz ; Christensen:2013va . Such a neutrino factory setup could be achieved with a proton beam with 2-4 MW Holtkamp:2000xn ; Ozaki:2001bb ; NeutrinoFactory:2004odt ; Geer:2006tb ; Ankenbrandt:2009zza .
@@ 5 Conclusions
We also investigated the role of charge identification of electrons or muons (or both). We found that CID can lead to some improvement in the sensitivity to CP violation and the other parameters. The impact is largest with larger statistics as CID helps to cut through systematics which become relevant at high statistics. We also found that perfect CID is not required to get the full benefit as significant enhancement occurs at potentially realistic LArTPC CID as well.
@@ Acknowledgements.
PBD acknowledges support by the United States Department of Energy under Grant Contract No. DE-SC0012704. JG acknowledges support by the U.S. Department of Energy Office of Science under award number DE-SC0025448.
@@ Appendix A Number of events
Figure 8: Number of events at a neutrino factory for the BNL-SURF and FNAL-SURF baseline with muon energy 8 GeV and 5 GeV for $10^{22}$ muon decays (nominally 10 years). We show the number of events from the $\nu_{e}\to\nu_{\mu}$ channel in red for varying values of $\delta$ and the dominant background from $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$.
@@
Appendix B Results for inverted ordering and higher energies for BNL setup
In tab. 3 we summarize the precision on the oscillation parameters for the BNL setup with higher muon energies of 12 GeV and 15 Gev.
In tab. 4 we show the precision for the oscillation parameters in inverted ordering.
Table 3: Results for precision on parameters for BNL baseline at higher energies. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays, nominally 10 years of each experiment.
$\delta^{\text{true}}=(-90^{\circ},~{}0)$ no CID 100% eCID 100% $\mu$CID
DUNE+HK+NF(BNL) E = 12 GeV $(10.8^{\circ},4.1^{\circ})$ $(7.9^{\circ},3.4^{\circ})$ $(8.6^{\circ},3.4^{\circ})$
DUNE+HK+NF(BNL) E = 15 GeV $(11.2^{\circ},4.2^{\circ})$ $(8.0^{\circ},3.7^{\circ})$ $(8.5^{\circ},3.6^{\circ})$
$\theta_{13}^{\text{true}}=8.54^{\circ}$
DUNE+HK+NF(BNL) E = 12 GeV $0.13^{\circ}$ $0.12^{\circ}$ $0.12^{\circ}$
DUNE+HK+NF(BNL) E = 15 GeV $0.13^{\circ}$ $0.12^{\circ}$ $0.12^{\circ}$
$(\Delta m_{31}^{2})^{\text{true}}=2.511\cdot 10^{-3}~{}\text{eV}^{2}$ $[10^{-6}\text{eV}^{2}]$ $[10^{-6}~{}\text{eV}^{2}]$ $[10^{-6}~{}\text{eV}^{2}]$
DUNE+HK+NF(BNL) E = 12 GeV 6.2 6.0 6.2
DUNE+HK+NF(BNL) E = 15 GeV $6.4$ $6.1$ 6.3
$\theta_{23}^{\text{true}}=(42^{\circ},~{}45^{\circ},~{}49^{\circ})$
DUNE+HK+NF(BNL) E = 12 GeV $(0.17^{\circ},0.67^{\circ},0.18^{\circ})$ $(0.16^{\circ},0.61^{\circ},0.18^{\circ})$ $(0.15^{\circ},0.61^{\circ},0.18^{\circ})$
DUNE+HK+NF(BNL) E = 15 GeV $(0.17^{\circ},0.73^{\circ},0.18^{\circ})$ $(0.16^{\circ},0.65^{\circ},0.18^{\circ})$ $(0.16^{\circ},0.67^{\circ},0.18^{\circ})$
Table 4: Resolution on the other parameters for combinations of different experiments assuming true inverted ordering. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays – nominally 10 years of each experiment unless otherwise specified. We also show the sensitivities for 960 kT-MW-yr of DUNE, twice its nominal exposure, combined with 2.47 MT-MW-yr of HK.
$\delta^{\text{true}}=(-90^{\circ},~{}0)$ no CID 100% eCID 100% $\mu$CID
HK $(20.2^{\circ},5.6^{\circ})$
DUNE $(18.8^{\circ},8.6^{\circ})$
DUNE+HK $(14.2^{\circ},4.7^{\circ})$
DUNE (20 yr)+HK $(11.4^{\circ},4.3^{\circ})$
DUNE+HK+NF(FNAL) $(12.4^{\circ},4.0^{\circ})$ $(9.7^{\circ},3.2^{\circ})$ $(9.8^{\circ},3.2^{\circ})$
DUNE+HK+NF(BNL) $(11.9^{\circ},4.2^{\circ})$ $(9.4^{\circ},3.5^{\circ})$ $(10.0^{\circ},3.6^{\circ})$
DUNE+HK+NF(BNL) $E=12$ GeV $(12.0^{\circ},4.3^{\circ})$ $(8.8^{\circ},3.7^{\circ})$ $(9.2^{\circ},3.7^{\circ})$
DUNE+HK+NF(BNL) $E=15$ GeV $(12.5^{\circ},4.4^{\circ})$ $(8.6^{\circ},3.9^{\circ})$ $(8.8^{\circ},3.8^{\circ})$
$\theta_{13}^{\text{true}}=8.54^{\circ}$
HK $0.22^{\circ}$
DUNE $0.22^{\circ}$
DUNE+HK $0.17^{\circ}$
DUNE (20 yr)+HK $0.15^{\circ}$
DUNE+HK+NF(FNAL) $0.16^{\circ}$ $0.14^{\circ}$ $0.13^{\circ}$
DUNE+HK+NF(BNL) $0.16^{\circ}$ $0.12^{\circ}$ $0.12^{\circ}$
DUNE+HK+NF(BNL) $E=12$ GeV $0.15^{\circ}$ $0.13^{\circ}$ $0.13^{\circ}$
DUNE+HK+NF(BNL) $E=15$ GeV $0.16^{\circ}$ $0.14^{\circ}$ $0.14^{\circ}$
$(\Delta m_{32}^{2})^{\text{true}}=-2.498\cdot 10^{-3}~{}\text{eV}^{2}$ $[10^{-6}\text{eV}^{2}]$ $[10^{-6}~{}\text{eV}^{2}]$ $[10^{-6}~{}\text{eV}^{2}]$
HK $8.9$
DUNE $8.9$
DUNE+HK $6.2$
DUNE (20 yr)+HK $5.1$
DUNE+HK+NF(FNAL) $5.0$ $4.8$ $5.0$
DUNE+HK+NF(BNL) $4.5$ $4.3$ $4.4$
DUNE+HK+NF(BNL) $E=12$ GeV $5.1$ $5.0$ $5.0$
DUNE+HK+NF(BNL) $E=15$ GeV $5.4$ $5.0$ $5.1$
$\theta_{23}^{\text{true}}=(42^{\circ},~{}45^{\circ},~{}49^{\circ})$
HK $(0.22^{\circ},1.33^{\circ},0.23^{\circ})$
DUNE $(0.35^{\circ},1.32^{\circ},0.35^{\circ})$
DUNE+HK $(0.18^{\circ},1.07^{\circ},0.19^{\circ})$
DUNE (20 yr)+HK $(0.16^{\circ},0.97^{\circ},0.17^{\circ})$
DUNE+HK+NF(FNAL) $(0.17^{\circ},1.01^{\circ},0.17^{\circ})$ $(0.17^{\circ},0.91^{\circ},0.17^{\circ})$ $(0.16^{\circ},0.86^{\circ},0.17^{\circ})$
DUNE+HK+NF(BNL) $(0.17^{\circ},0.81^{\circ},0.17^{\circ})$ $(0.16^{\circ},0.68^{\circ},0.17^{\circ})$ $(0.15^{\circ},0.70^{\circ},0.16^{\circ})$
DUNE+HK+NF(BNL) $E=12$ GeV $(0.17^{\circ},0.97^{\circ},0.17^{\circ})$ $(0.17^{\circ},0.78^{\circ},0.17^{\circ})$ $(0.17^{\circ},0.86^{\circ},0.18^{\circ})$
DUNE+HK+NF(BNL) $E=15$ GeV $(0.18^{\circ},1.00^{\circ},0.18^{\circ})$ $(0.17^{\circ},0.84^{\circ},0.18^{\circ})$ $(0.17^{\circ},0.91^{\circ},0.18^{\circ})$